OpenStudy (anonymous):

What is the integral of (sqrt(1 - x^2) / x)? Thanks

OpenStudy (anonymous):

let x = sinθ, then: sqrt(1 - x^2) / x = sqrt(1 - (sinθ)^2)/sinθ = sqrt((cosθ)^2)/sinθ = cosθ/sinθ = cotθ then take the integral of cotθ

OpenStudy (cruffo):

you will also have to replace \(dx\) with \(\cos\theta d\theta\)

OpenStudy (cruffo):

So when you let \(x =\sin\theta\) then \(dx = \cos\theta d\theta\). Then \[\int \frac{\sqrt{1-x^2}}{x} dx = \int \frac{\cos^2\theta }{\sin\theta} d\theta\] Rewrite \(\cos^2\theta\) as \(1- \sin^2\theta\) and break up the fraction, \[\int \frac{\cos^2\theta }{\sin\theta} d\theta = \int \frac{1 - \sin^2\theta }{\sin\theta} d\theta = \int (csc\theta - \sin\theta) d\theta\]

OpenStudy (anonymous):

Thanks!